### v2

# Training-free Anomaly Event Detection via LLM-guided Symbolic Pattern Discovery

Abstract
Anomaly event detection plays a crucial role in various real-world applications. However, current approaches predominantly rely on supervised learning, which faces significant challenges: the requirement for extensive labeled training data and lack of interpretability in decision-making processes. To address these limitations, we present a training-free framework that integrates open-set object detection with symbolic regression, powered by Large Language Models (LLMs) for efficient symbolic pattern discovery. The LLMs guide the symbolic reasoning process, establishing logical relationships between detected entities. Through extensive experiments across multiple domains, our framework demonstrates several key advantages: (i) achieving superior detection accuracy through direct reasoning without any training process; (ii) providing highly interpretable logical expressions that are readily comprehensible to humans; and (iii) requiring minimal annotation effort - approximately 1% of the data needed by traditional training-based methods. To facilitate comprehensive evaluation and future research, we introduce two datasets: a large-scale private dataset containing over 110,000 annotated images covering various anomaly scenarios including construction site safety violations, illegal fishing activities, and industrial hazards, along with a public benchmark dataset of 5,000 samples with detailed anomaly event annotations. Code is available at here.
1 Introduction
Visual Anomaly Detection (VAD) serves as a cornerstone technology in modern computer vision systems, focusing on identifying irregular patterns or unexpected deviations from established normality in visual data. This field has witnessed exponential growth in its practical applications, playing critical roles in industrial quality control [51, 58], medical diagnosis [54], surveillance systems [32], and autonomous driving safety [5]. The increasing demand for robust anomaly detection solutions has sparked intensive research efforts, leading to significant advancements in bridging the gap between academic innovations and industrial deployments [36, 2, 3, 8, 13, 62, 65, 29]. Despite these advances, developing efficient and reliable VAD systems remains challenging, particularly in scenarios requiring rapid adaptation to new anomaly patterns while maintaining interpretable decision processes.
Figure 1: Illustration of training-based methods alongside out VED-SR pipeline. VED-SR provides low-consumption and plug-and-play solutions.
Despite considerable advancements in Visual Anomaly Detection (VAD), Event-based Anomaly Detection (EAD) involving complex object interactions still presents substantial challenges that have not been fully addressed. There are three main limitations in current methods. First, these methods largely depend on extensive annotated training datasets, which are often limited or extremely costly to acquire in practical settings [19]. This reliance restricts their applicability in various domains. Second, current techniques struggle to grasp the intricate relationships among entities [7]. The significant semantic differences between normal and abnormal behaviors, along with a lack of prior knowledge about anomalies, make it difficult to understand and formulate complex logical rules, particularly in situations with numerous interacting entities or different quantities. Third, the environments in which event detection systems operate, such as processing vast data streams from surveillance cameras and unmanned aerial vehicles [20], present notable scalability issues. The infrequency of anomalies and the marked imbalance between normal and abnormal activities in real-world data further complicate the detection process, rendering traditional training-based methods increasingly ineffective.
Figure 2: The overall performance on the Multi-Event benchmark. Our method outperforms other multimodal fine-tuning methods under multiple abnormal events without requiring training.
To address these challenges, existing approaches can be broadly categorized into two paradigms: implicit modeling and explicit relationship analysis. The implicit modeling approaches aim to capture inter-entity relationships through end-to-end feature learning. GCAD [4] pioneered this direction by employing dual student-teacher architectures for comprehensive anomaly detection. Subsequent works have enhanced this framework: DSKD [67] introduced contextual affinity distillation to better understand normal global context, while GLCF [61] adopted a local-global alignment strategy with a bottleneck architecture. More recently, EfficientAD [1] achieved computational efficiency through knowledge distillation while maintaining relationship modeling capability. The explicit relationship analysis approaches, on the other hand, adopt a more interpretable two-stage strategy: first detecting individual entities, then analyzing their relationships. ComAD [31] exemplifies this direction by combining unsupervised semantic segmentation through K-means clustering on DINO pre-trained features with metrological feature analysis, offering a training-free solution. PSAD [23]further advances this paradigm through few-shot learning, integrating positional features and multiple memory banks for modeling component quantities, arrangements, and fine-grained features. However, while both paradigms have shown promise, they still face fundamental limitations: implicit methods lack interpretability and require substantial training data, while existing explicit methods struggle with complex logical relationships and computational efficiency.
To overcome these limitations, we propose Visual Event Detection with Symbolic Reasoning (VED-SR), a novel training-free framework that synergizes symbolic regression with Large Language Model (LLM) reasoning for anomaly event detection. Our approach fundamentally differs from existing solutions in three key aspects, details in Figure 1. First, rather than relying on direct pattern matching or extensive training, VED-SR leverages LLMs’ symbolic reasoning capabilities and rich commonsense knowledge to construct contextually appropriate anomaly detection models. This enables more nuanced and adaptable anomaly pattern recognition across diverse scenarios. Second, we incorporate symbolic regression, a classical interpretable modeling approach that has demonstrated remarkable effectiveness in capturing intrinsic data patterns while maintaining strong generalization capabilities . The integration of symbolic regression with LLM guidance creates a powerful synergy: the LLM’s prior knowledge steers the symbolic pattern discovery process, while symbolic regression ensures interpretable and generalizable results.
Our framework adopts a training-free, plug-and-play architecture that makes it immediately deployable across different domains without requiring domain-specific training. Extensive experiments across multiple visual event benchmarks, including UCSD Ped2 [52], ERA [35], and our newly introduced Multi-Event and Helmet-Mac datasets, demonstrate the effectiveness of VED-SR. Notably, when integrated with three distinct open-source multimodal object detectors, our method achieves an AUROC of 98.7% on the UCSD Ped2 dataset without any fine-tuning, approaching the state-of-the-art performance of 99.7% achieved by training-based methods. Furthermore, as illustrated in Figure 2, our approach demonstrates comparable and occasionally leading performance on the Multi-Event Dataset. These results, combined with consistent superior performance across other benchmarks, validate the effectiveness and generalizability of our approach. To summarize, Our key contributions include:
  • We propose VED-SR, a novel training-free framework for anomaly event detection that integrates open-set object detection with symbolic regression, guided by Large Language Models. This approach eliminates the need for training data while maintaining competitive detection performance.
  • We develop an efficient symbolic pattern discovery mechanism leveraging LLMs’ reasoning capabilities, which enables the automatic construction of interpretable logical expressions for anomaly detection. This mechanism significantly reduces the annotation effort to approximately 1% of what traditional training-based methods require.
  • We introduce two comprehensive anomaly event detection datasets: a large-scale private dataset containing over 110,000 annotated images covering diverse scenarios (construction site safety violations, illegal fishing activities, and industrial hazards), and a public benchmark dataset of 5,000 samples with detailed annotations to facilitate reproducible research.
  • Extensive experiments demonstrate that our training-free approach achieves comparable or superior performance to state-of-the-art training-based methods across multiple benchmarks (e.g. achieving 98.7% AUROC on the UCSD Ped2 dataset without fine-tuning while demonstrating a remarkable 15% performance improvement on the Helmet-Mac dataset), while providing explicit reasoning paths for human verification.
2 Related work
2.1 Visual Anomaly Detection
Visual anomaly detection [59] has emerged as a fundamental research direction in computer vision, with various approaches developed to identify deviations from normal patterns in visual data. These approaches can be broadly categorized into several paradigms: memory-based, reconstruction-based, and more recently, vision-language model-based methods. Memory-based approaches, exemplified by PatchCore [41], directly store and utilize normal sample features for anomaly scoring through nearest-neighbor calculations. Reconstruction-based methods, such as DFR [47], employ auto-encoders to rebuild normal features, while knowledge distillation approaches like RD4AD [11] and ViTAD [66] leverage student networks for feature regression. To address the challenge of over-generalization in anomaly scoring, methods like Draem [63] and MRKD [21] introduce synthetic anomalies to enhance model discrimination capabilities. Recent advancements have explored novel directions. RegAD [18] introduces registration-based proxy tasks for representation learning, while GraphCore [58] employs vision isometric invariant graph neural networks for structural feature extraction. The emergence of vision-language models has opened new possibilities, with WinCLIP [19] pioneering the use of CLIP for anomaly detection through text-image similarity computation. This direction has been further enhanced by APRIL-GAN [9] and AnomalyCLIP [68], which address the limitations of manual prompt design and improve model adaptability. Despite these advances [2, 3, 52, 5, 42], the specific domain of anomaly event detection lacks comprehensive datasets and effective approaches for handling complex scenarios. Our work addresses this gap by introducing a dataset focused on safety helmet compliance in construction sites, characterized by complex backgrounds and long-tail distributions of anomalous events
2.2 Prompt Learning
Prompt Learning represents a paradigm shift in utilizing pre-trained models, fundamentally transforming the approach to downstream tasks [53]. Unlike traditional fine-tuning methods, prompt learning aligns downstream tasks with pre-trained models’ inherent capabilities, demonstrating remarkable versatility across various applications. The field’s evolution began with the LAMA dataset [38], which demonstrated the effectiveness of manually crafted cloze templates in probing language model knowledge. Subsequent research has shown that engineered prefix prompts can effectively handle diverse tasks [6], ranging from basic operations to complex reasoning challenges. In the vision-language domain, prompt-based approaches have become increasingly prevalent in anomaly detection applications [43, 45, 44, 48]. Our work extends this paradigm by designing a specialized prompt space for anomaly event detection, leveraging LLMs’ symbolic reasoning capabilities and commonsense knowledge to guide symbolic pattern discovery.
2.3 Open-set Object Detection
The integration of large language models and multimodal learning frameworks has revolutionized open-set object detection [24, 27, 55, 56, 57, 60, 64, 69, 70, 71, 72]. MDETR [22] pioneered this evolution by introducing text-guided visual region alignment through a DETR-like architecture. GenerateU [28] advanced the field by enabling object identification and naming without predefined categories, while GLIP [26] and GroundingDINO [30] demonstrated the effectiveness of combining detection with language grounding. Recent developments include YOLO-World [10], which achieves efficient zero-shot detection through region-text pair pre-training, and APE [46], which unifies detection and grounding capabilities. While these advances have significantly expanded the capabilities of open-set detection, their application to anomaly event detection remained unexplored until our work with VED-SR.
3 Method
Figure 3: Illustration of the proposed VED-SR. VED-SR mainly consists of logic search and symbolic reasoning. The former module constructs and explores the search space by leveraging structured entity features extracted from an open-set object detector. Th latter module harnesses the symbolic reasoning capabilities of Large Language Models (LLMs) along with their inherent commonsense understanding of anomaly patterns to guide the search process toward more appropriate and rational pathways.
Recent advances in anomaly detection have predominantly focused on developing lightweight and zero-shot detection frameworks. However, achieving training-free operation while maintaining interpretability remains a significant challenge in anomaly event detection. To address this gap, we propose VED-SR, a novel framework that synergistically combines symbolic regression with Large Language Model (LLM) reasoning. Our approach consists of two key components: (1) a symbolic logic search module that discovers interpretable patterns through entity-level analysis, and (2) an automated LLM reasoning module that leverages semantic understanding to guide and accelerate the pattern discovery process. Figure 3 illustrates our framework’s architecture. Formally, given a visual dataset D containing potential anomalous events:
$D=\{x^{n_{1}},x^{n_{2}},...,x^{n_{i}},x^{a_{1}},x^{a_{2}},...,x^{a_{j}}\}$ (1)
where $x^{n_{i}}$ and $x^{a_{j}}$ represent normal and abnormal samples respectively, our goal is to discover a discriminative function $f(\tau)$ that maps input samples to binary anomaly decisions:
$f(\tau):\mathcal{X}\rightarrow\{0,1\}$ (2)
where $\tau$ represents the parameters of the function. Unlike traditional approaches that learn $f(\tau)$ through supervised training, our method discovers interpretable symbolic expressions through an LLM-guided search process. In the following sections, we detail each component of our framework and explain how they work together to achieve efficient and interpretable anomaly detection.
3.1 Symbolic Logic Search
The core of our anomaly detection framework lies in discovering interpretable symbolic patterns that characterize anomalous events. This section details our symbolic logic search mechanism, which consists of entity extraction, feature representation, and symbolic pattern discovery.
Entity-level Feature Extraction. We first leverage an open-vocabulary object detector to extract comprehensive entity information from each sample. For a given image $x$, we obtain a set of entities:
$E=\{e_{1},e_{2},...,e_{n}\},\quad e_{i}=(c_{i},b_{i},s_{i})$ (3)
where $c_{i}$ represents the category label, $b_{i}=(x,y,w,h)$ denotes the bounding box coordinates, and $s_{i}$ is the detection confidence score. These entities form the basis for our symbolic pattern analysis. To facilitate symbolic reasoning, we transform the raw entity information into structured features:
$\mathbf{X}=\{\phi_{1}(E),\phi_{2}(E),...,\phi_{d}(E)\}$ (4)
where $\phi_{i}(\cdot)$ represents different feature extraction functions that capture entity counts, spatial relationships, and attribute distributions.
Symbolic Pattern Discovery Given the extracted features, we employ symbolic regression to discover interpretable patterns that distinguish normal from anomalous events. The symbolic regression problem is formulated as:
$f^{*}=\arg\min_{f\in\mathcal{F}}\sum_{i=1}^{n}\mathcal{L}(f(\mathbf{X}_{i}),y_{i})+\lambda\Omega(f)$ (5)
[…]
  1. 1.
    Initializes a population of candidate expressions
  2. 2.
    Evaluates their fitness on the current dataset
  3. 3.
    Applies genetic operators (mutation, crossover) to generate new candidates
  4. 4.
    Selects the best expressions for the next generation
This process generates human-interpretable symbolic expressions that capture meaningful patterns in the data. For example, in a safety helmet detection scenario, a discovered pattern might be:
$f(\mathbf{X})=\bigvee_{i\in\{p,d\}}[\phi_{i}(E)>\phi_{h}(E)]$ (6)
where $\phi_{p}$, $\phi_{h}$, and $\phi_{d}$ represent the counting functions for persons, helmets, and heads respectively. The discovered patterns serve as initial hypotheses that will be further refined through LLM reasoning, as detailed in the next section.
3.2 Automated LLM Reasoning
To enhance the efficiency and effectiveness of symbolic pattern discovery, we propose an automated reasoning mechanism that leverages the semantic understanding capabilities of Large Language Models (LLMs). This section details our structured prompt space design and its integration with symbolic logic search.
3.2.1 Structured Prompt Space
We design a hierarchical prompt space to facilitate effective communication with LLMs through three key components:
Scene Context Initialization. The first layer of prompts establishes the scene context:
$P_{init}=\{\text{scene},\text{entities},\text{constraints}\}$ (7)
This activation prompt triggers the LLM’s prior knowledge relevant to the specific anomaly detection scenario, creating crucial connections between visual entities and semantic understanding.
Chain-of-Thought Guidance. The second layer provides structured reasoning steps:
$P_{cot}=\{s_{1}\rightarrow s_{2}\rightarrow...\rightarrow s_{k}\}$ (8)
where each step $s_{i}$ guides the LLM through professional analytical frameworks for identifying potential anomaly patterns. This systematic approach ensures comprehensive consideration of entity relationships and domain constraints.
Contextual Feedback Integration. The final layer incorporates evaluation feedback:
$P_{feed}=\{(r_{1},\alpha_{1}),(r_{2},\alpha_{2}),...,(r_{n},\alpha_{n})\}$ (9)
where $r_{i}$ represents previous reasoning attempts and $\alpha_{i}$ their corresponding effectiveness scores. This feedback mechanism enables the LLM to refine its suggestions based on historical performance.
Algorithm 1 LLM-Guided Symbolic Search
1:  Initialize population $\mathcal{P}_{0}$ of symbolic expressions
2:  for each iteration $t$ do
3:     $f_{t}^{*}\leftarrow$ Select best expression from $\mathcal{P}_{t}$
4:     $S_{t}\leftarrow$ Generate LLM suggestions via prompt space
5:     $\mathcal{P}_{t+1}\leftarrow$ Update population using $\{f_{t}^{*},S_{t}\}$
6:     if convergence criterion met then
7:        break
8:     end if
9:  end for
10:  return  best expression $f^{*}$
3.2.2 LLM-Guided Symbolic Search
We establish a bidirectional interaction mechanism between LLM reasoning and symbolic logic search (SLS) through an iterative process: At each iteration, the best symbolic expression $f_{t}^{*}$ serves as a directional indicator for LLM reasoning. The LLM analyzes this expression through our structured prompt space and generates suggestions $S_{t}$ that incorporate its semantic understanding and common sense knowledge. These suggestions are then transformed into new symbolic expressions and integrated into the population for the next iteration. The integration creates a synergistic effect where:
  • LLM reasoning guides the symbolic search towards semantically meaningful patterns
  • Symbolic search provides objective evaluation of LLM suggestions
  • The iterative process combines the interpretability of symbolic expressions with the rich semantic understanding of LLMs
This bidirectional interaction accelerates the discovery of meaningful anomaly patterns while maintaining the interpretability of the detection process. The LLM’s suggestions help navigate the vast space of possible symbolic expressions, while the symbolic search framework ensures that the final patterns remain explicit and verifiable.
Analysis. Through the integration of symbolic logic search and LLM reasoning, our approach offers several distinct advantages. First, the training-free nature of our framework eliminates the need for extensive labeled datasets, as the symbolic pattern discovery process operates directly on entity-level features extracted from test samples. This characteristic is particularly valuable in real-world scenarios where collecting and annotating training data is costly or impractical. Second, our method achieves high interpretability through two complementary mechanisms: (a) the discovered symbolic expressions provide explicit, human-readable logical patterns that directly explain the detection decisions, and (b) the LLM’s reasoning process offers semantic context for these patterns, making them more accessible to domain experts. This dual-layer interpretability sets our approach apart from traditional black-box models. Third, the bidirectional interaction between symbolic search and LLM reasoning creates an efficient optimization process. The LLM’s semantic understanding and common sense knowledge help navigate the vast space of possible symbolic expressions, significantly reducing the search complexity compared to pure evolutionary approaches. Meanwhile, the symbolic search framework provides objective evaluation criteria that prevent the accumulation of LLM reasoning errors, ensuring the robustness of the final patterns. Lastly, our structured prompt space design enables systematic utilization of LLM capabilities while maintaining consistency in the reasoning process. This structured approach not only improves the quality of LLM suggestions but also makes the entire framework more stable and reproducible compared to methods relying on ad-hoc prompting strategies.
4 Experiment
Table 1: Performance of different open set detectors on multiple data sets with or without VED-SR module. (AUROC%)
Datasets APE [46] YOLO-World [10] GLIP [26]
Original +VED-SR Original +VED-SR Original +VED-SR
BALL 55.36 94.91 (+39.55) 54.76 89.05 (+34.29) 66.34 90.27 (+23.93)
ERA [35] PersonCrowd 78.30 83.26 (+4.96) 55.00 85.11 (+30.11) 81.71 85.08 (+3.37)
Sport 67.13 90.29 (+23.16) 67.27 88.54 (+21.27) 66.94 89.65 (+22.71)
Helmet-Mac 67.41 83.18 (+15.77) 65.40 82.47 (+17.07) 61.06 76.25 (+15.19)
Multi-rods Fishing1 66.82 75.16 (+8.36) 52.72 72.01 (+19.29) 50.00 71.11 (+21.11)
  • 1
    It refers to a subset of Multi-Event Dataset.
4.1 Experimental Setup
To comprehensively evaluate our approach, we conduct experiments on four diverse datasets: two large-scale self-collected datasets and two public benchmarks, each targeting different aspects of anomaly event detection.
4.1.1 Self-collected Datasets
Multi-Event Dataset is our large-scale dataset about detecting multiple anomaly event in complex scenarios. The dataset comprises over 110,000 images spanning various scenarios including fires and waste incineration (30,954 images), multi-rods fishing (12,000 images), night fishing (97 images), license plate detection (12,487 images), personnel loitering and intrusion (10,788 images), among others. For this study, we specifically focus on the multiple-rod fishing scenario, where we have 15,000 training images with 45,341 detailed bounding box annotations covering persons, fishing rods, tackle bags, and umbrellas. The test set contains 2,283 images, with 1,098 cases showing anomalous multi-rods fishing activities and 1,185 normal cases with single-rod fishing. The rich annotations enable precise analysis of entity relationships and logical constraints.
Helmet-Mac Dataset, which we make publicly available [^1] Dataset will soon be available in our code repository., addresses the critical domain of construction site safety monitoring. This dataset is curated from various construction scenarios and focuses on safety helmet compliance detection. It contains 7,571 training images with detailed annotations of human heads and safety helmets across diverse construction environments. The test set comprises 4,642 images, balanced between 2,276 safety violations (workers without helmets) and 2,366 compliant cases. The dataset captures various challenging scenarios including different lighting conditions, viewing angles, and occlusion cases, making it a valuable benchmark for safety-critical event detection systems.
4.1.2 Public Benchmarks
ERA Dataset [35] (Event Recognition in Aerial videos) provides a comprehensive collection of aerial footage covering various event categories. We organize our evaluation around three main event categories: BALL events (327 images) encompassing baseball, soccer, and basketball games; Person_crowded events (352 images) including concerts, conflicts, parade protests, and parties; and Sport events (258 images) covering cycling, boating, and racing activities. Additionally, we utilize 347 Non-event images as negative samples, creating a balanced evaluation framework for our method’s discriminative capabilities across different event types.
UCSD Ped2 Dataset [52] serves as our primary benchmark for comparison with state-of-the-art methods. This well-established dataset has been widely used in the anomaly detection community, providing a standardized evaluation platform. We use this dataset to demonstrate our method’s competitive performance against existing approaches while maintaining the advantages of training-free operation and interpretability.
4.1.3 Implementation Details
Our implementation consists of three main components: multimodal detection, symbolic regression, and computational infrastructure. We detail each aspect to ensure reproducibility.
Multimodal Detection Setup We employ three state-of-the-art multimodal detectors in our experiments: APE [46] (serving as the primary detector), GLIP [26], and YOLO-WORLD [10]. To optimize detection performance, we implement a two-stage prompt generation process. Initially, we leverage LLM to analyze event scenario descriptions and generate comprehensive detection prompts. The LLM not only generates prompts for objects directly related to the event scenarios but also produces prompts for contextually related but non-event objects, ensuring broad coverage of potential scene elements. Based on empirical studies of detector characteristics, we configure different detection thresholds for optimal performance. Considering APE’s characteristically lower threshold nature, we set its minimum detection threshold to 0.05. For GLIP and YOLO-WORLD, we establish a higher threshold of 0.1 to maintain a balance between precision and recall in object detection.
Symbolic Regression Configuration The symbolic regression module processes the detection results through an iterative optimization procedure. Upon receiving detection outputs, the module generates initial logical expressions and evaluates their fitness. If termination criteria are not met, it selects the top-4 logical expressions for crossover mutation, continuing this process until reaching optimal expressions. We configure the symbolic regression parameters based on extensive experimental validation. The population size is set to twice the number of target categories, allowing for sufficient expression diversity. The crossover and mutation factors are set to 0.5 and 0.3 respectively, providing a balanced exploration-exploitation trade-off. The optimization process continues for 5,000 iterations or until convergence criteria are satisfied.
Computational Resources Our experimental framework is deployed on a cluster of 8 NVIDIA V100 GPUs, each equipped with 32GB memory. This computational infrastructure supports three main tasks: the inference operations of base object detectors, the deployment and prompt processing of Large Language Models, and the computational requirements of symbolic regression optimization. This hardware configuration ensures efficient processing of our training-free anomaly detection pipeline while maintaining real-time performance capabilities.
Table 2: The overall performance on the UCSD ped2 [52] benchmark.This intuitively reflects that the performance of VED-SR (APE) without training is very close to the current SOTA.
Training-free Methods score (%)
ConvVQ [33] 90.2
SD-MAE [40] 95.4
FastAno [37] 99.3
$\times$ VALD-GAN [49] 97.74
MAMA [16] 98.2
Backgroud-Agnostic [15] 98.7
DMAD [33] 99.7
VED-SR 98.7
4.2 Main Results
We conduct comprehensive experiments to evaluate our approach from multiple perspectives: effectiveness of the training-free framework, necessity of fine-tuning, and comparison with traditional methods.
Table 3: Performance of different fine-tuned methods and with or without VED-SR module. Lora indicates whether to Lora-tuning the APE model. Prompt indicates whether to Prompt-tuning the APE model (AUROC%)
VED-SR Lora Prompt Helmet-Mac Multi-rods-Fishing
67.41 66.82
83.18 (+15.77) 75.16 (+8.34)
84.11 75.86
95.67 (+11.56) 78.44 (+2.58)
67.42 66.82
81.62 (+14.2) 76.06 (+9.24)
Comparison with Different Detection Architectures. We first evaluate our framework using three state-of-the-art open-vocabulary detectors: APE, GLIP, and YOLO-WORLD, which represent diverse architectural choices in both visual and language processing. These detectors employ different visual backbones (VIT [14], Swin-L [34], and YOLOv8) and language models (EVA-CLIP [50], CLIP [39], and BERT [12]). As shown in Table 1, all three detectors achieve strong performance without any fine-tuning, with APE consistently outperforming others across all five anomaly event scenarios. This superior performance of APE can be attributed to its more sophisticated visual-language alignment mechanism and larger pre-training dataset, which enables better transfer of knowledge to anomaly detection tasks. The consistent performance across architecturally diverse models also suggests that our framework’s effectiveness is not tied to specific architectural choices, but rather stems from the fundamental synergy between symbolic reasoning and detection capabilities. Finding 1: The effectiveness of our training-free framework is architecture-agnostic, with APE’s superior performance likely due to its enhanced visual-language alignment and broader pre-training.
Comparison with Fine-tuning Approaches. To further validate the efficiency of our training-free approach, we conduct comparative experiments with fine-tuned variants on the Helmet-Mac and Multi-rods Fishing datasets. We implement two common fine-tuning strategies: LORA fine-tuning[17] and Prompt tuning [25], representing different levels of parameter adaptation. Results in Table 3 demonstrate that our training-free approach achieves comparable performance to fine-tuned models. This intriguing finding suggests that our training-free approach effectively leverages the model’s general understanding of visual-language relationships. Additionally, the symbolic reasoning component provides a more structured way to capture anomaly patterns compared to implicit learning through fine-tuning, achieving similar effectiveness without the computational overhead of parameter adaptation. Finding 2: Our training-free approach is comparable to fine-tuning methods, possibly due to better preservation of general visual-language understanding and more structured pattern discovery.
Comparison with Traditional Methods. For comparison with traditional anomaly detection methods, we evaluate our approach on UCSD Ped2 [52] dataset. Our method achieves an impressive 98.7% accuracy without utilizing semantic-level annotations or task-specific fine-tuning, approaching the state-of-the-art performance (99.7%), showed in Table 2. This minimal performance gap is particularly interesting considering the vast difference in approach complexity. We hypothesize that this effectiveness stems from two factors: first, the pre-trained detectors already possess rich semantic understanding that generalizes well to anomaly detection; second, our symbolic reasoning framework effectively translates this semantic knowledge into explicit detection rules, potentially capturing patterns that are similar to those learned by supervised methods but in a more interpretable manner. Finding 3: The near-SOTA performance on UCSD Ped2 suggests that combining pre-trained knowledge with symbolic reasoning can effectively match supervised learning capabilities.
Figure 4: Performance on VED-SR with or without LLM.
4.3 Ablation Study
Analysis of Component Contributions. To understand the contribution of each component in our framework, we conduct comprehensive ablation studies examining the individual and combined effects of LLM reasoning and symbolic regression. Starting with a baseline using only manual logic expressions (67.00% average performance), the addition of symbolic regression significantly improves performance to 85.36%. This substantial improvement (+18.36%) suggests that automated pattern discovery through symbolic regression is significantly more effective than human-designed rules, likely due to its ability to explore a broader space of logical combinations and capture subtle patterns that might not be immediately apparent to human experts.
Impact of LLM Integration. Figure 4 shows further integration of LLM reasoning not only enhances detection accuracy but accelerates the convergence of symbolic regression, showing a synergistic effect between semantic understanding and pattern discovery. A detailed analysis of the convergence process reveals that LLM guidance helps avoid many unnecessary exploration paths in the symbolic search space, effectively pruning unpromising directions based on semantic understanding. Moreover, we observe the quality of discovered patterns improves with LLM guidance, as the patterns tend to be more semantically meaningful and better aligned with human intuition about anomalies. Finding 4: The significant performance gains from each component suggest that effective anomaly detection requires both systematic pattern discovery (through symbolic regression) and semantic guidance (through LLM reasoning).
These ablation results highlight an important insight: while traditional approaches often treat anomaly detection as either a purely statistical or purely semantic task, our results suggest that combining both aspects leads to superior performance. The symbolic regression provides the systematic exploration capability, while LLM reasoning contributes semantic understanding, creating a framework that leverages the strengths of both approaches. Our experiments demonstrate that the combination of training-free operation, symbolic reasoning, and LLM guidance creates a robust and effective framework for anomaly detection, often matching or exceeding the performance of more complex, training-dependent approaches.
5 Conclusion
In this paper, we present VED-SR, a novel approach that addresses visual event detection through symbolic logical reasoning. Our framework makes three key contributions to the field: First, we establish a new paradigm by uniquely combining symbolic regression for logical pattern search with Large Language Models (LLMs) for symbolic reasoning. This integration successfully transforms complex anomaly event patterns into human-readable symbolic expressions while achieving performance comparable to current state-of-the-art methods. Second, we demonstrate the effectiveness of a training-free framework that eliminates the need for task-specific fine-tuning while maintaining competitive performance. This achievement suggests a promising direction for developing more efficient and adaptable event detection systems. Third, we introduce a new benchmark dataset containing over 5,000 annotated samples for visual event detection, providing a valuable resource for future research in this domain. Looking forward, several promising directions emerge for extending this work, including expanding the symbolic logic search space, automating prompt space construction, and integrating with object detection and active learning frameworks. These developments could further advance the field of visual event detection while maintaining the interpretability and efficiency advantages demonstrated in our work.